Using Table 2.13 below, if the expected time for activity E is changed to nine weeks, by how many weeks will the project's expected completion time increase?
Table 2.13
(All activity times for the project are in weeks.)

Activity

Optimistic Time

Most Likely Time

Pessimistic Time

Immediate Predecessors(s)

A

3

4

5

---

B

5

7

9

---

C

2

3

10

A

D

2

5

8

B

E

1

7

7

B

F

3

5

7

CD

G

7

8

9

DE

H

2

6

10

F
Answer #1

Critical path = Longest path duration of project = BDFH = 23 weeks

Activity ACTIVITY before Optimistic (a) Most likely (M) Pessimistic (b) PERT Estimate = (a+4*M+b)/6
A --- 3 4 5 4.00
B --- 5 7 9 7.00
C A 2 3 10 4.00
D B 2 5 8 5.00
E B 1 7 7 6.00
F CD 3 5 7 5.00
G DE 7 8 9 8.00
H F 2 6 10 6.00

New Expected time for E = 9 weeks. The critical path shifts to BEG. New project duration = 7 + 9+ 8 = 24 weeks